3. Carbaugh Corporation has provided the following production and average cost data for two levels of...

3.

Carbaugh Corporation has provided the following production and average cost data for two levels of monthly production volume. The company produces a single product.

  Production volume 5,900 units 6,900 units
  Direct materials $76.90 per unit $76.90 per unit
  Direct labor $51.20 per unit $51.20 per unit
  Manufacturing overhead $76.40 per unit $66.40 per unit

The best estimate of the total cost to manufacture 6,200 units is closest to: (Do not round intermediate calculations.)
Answer #1

Direct materials and direct labor are variable costs as the cost per output is same.

Total manufacturing overhead for 5900 units = 5900*76.40 = 450,760

Total manufacturing overhead for 6900 units = 6900*66.40 = 458,160

Variable Manufacturing Overhead = (Cost at highest activity - Cost at lowest activity)/(Highest activity-Lowest activity)

= (458,160-450,760)/(6900-5900) = 7.4

Fixed manufacturing overhead = 458,160 - (6900*7.4) = 407,100

.

Total Cost for 6,200 units

= Variable cost (76.90+51.20+7.40)*6200 +Fixed cost 407,100

= 1,247,200